If logxlog18 (Sq.rt 2 + Sq.rt 8) = 1/3. Then what is the value of 1000x?

Grade:11

1 Answers

Arun
4 years ago
logx (log18(√2+√8)) = 1/3
 
x1/3 = log18(√2+√8)
 
x1/3 = log18(3√2)
 
     √18 = √9√2 = 3√2
 
So, x1/3 = log18(18)1/2
 
      x1/3 = 1/2
 
      (x1/3)3 = (1/2)3
 
           x = 1/8. So, 1000x = 125.